OpenStudy (anonymous):

Write a set of linear inequalities for the production of boots and jackets if there are 800 hours available for making both items and 1200 hours for finishing both items and a pair of boots takes 4 hours to make and 3 hours to finish and a jacket takes 2 hours to make and 4 hours to finish. Market experience states that more than 150 boots have to be produced every month.

OpenStudy (anonymous):

My answers are\[4x + 2y \le 800 , 3x + 4y \le 1200 , x \ge 150\]

OpenStudy (anonymous):

are these right or am I being to simple?

ganeshie8 (ganeshie8):

looks good, except for the last inequality

OpenStudy (anonymous):

the x has to be greater than or equal to 150?

ganeshie8 (ganeshie8):

Market experience states that more than 150 boots have to be produced every month that gives you \(x \gt 150\)

OpenStudy (anonymous):

Ahh ok not equal to then.

ganeshie8 (ganeshie8):

yes, it must be strictly greater than 150. if the q says, atleast 150 booths, then we wud have equal sign also..

ganeshie8 (ganeshie8):

*boots

OpenStudy (anonymous):

Thank you for that clarification.

ganeshie8 (ganeshie8):

np :)